### v1

@@ 3.1 Overlapping layer-wise additive Schwarz preconditioner
The computational cost associated with solving the preconditioned system of equations (10) is directly linked to the quality of the nonlinear preconditioner $\pazocal{M}$. In this work, we construct the preconditioner $\pazocal{M}$ by utilizing the overlapping layer-wise network decomposition. Thus, we decompose the DNN, consisting of $N$ layers, into $N_{sd}$ subnetworks (subdomains), such that each layer belongs to a unique subnetwork. The parameters of the network are also decomposed into $N_{sd}$ disjoint groups, i.e.,
$\boldsymbol{\theta}=[\boldsymbol{\theta}_{1},\cdots,\boldsymbol{\theta}_{s},\cdots,\boldsymbol{\theta}_{N_{sd}}]^{T},$
where $\boldsymbol{\theta}_{s}:=[\boldsymbol{\theta}_{s,1},\ldots,\boldsymbol{\theta}_{s,N_{s}}]^{T}\in\mathbb{R}^{n_{s}}$, for $s=1,\ldots,N_{sd}$. Each $s^{th}$ non-overlapping subnetwork consists of $N_{s}$ layers, with parameters of $h^{th}$ layer being denoted as $\boldsymbol{\theta}_{s,h}$. In case of the overlapping decomposition, we concatenate parameters $\boldsymbol{\theta}_{s}$ with the parameters of neighboring $\delta(\in\mathbb{N})<N_{S}$ layers from the $(s-1)$-th and $(s+1)$-th subnetworks. For $\delta\geq 1$, this gives rise to overlapping set of subdomain parameters
@@
where $\tilde{N}_{s}$ and $\tilde{n}_{s}$ denotes the number of layers in overlapping subnetwork and the number of associated subnetwork parameters, respectivelly. Note that, for the first ($s=1$) and the last ($s=N_{sd}$-th) subnetworks, we only pick layers from the second and $(N_{sd}-1)$-th subnetworks, respectively. Fig. 2 illustrates the parameter selection for the second subnetwork of the six-layer network decomposed in a layer-wise manner into three subdomains with an overlap of size one, i.e., $N_{sd}=3$ and $\delta=1$.
@@ 3.1.1 Single-level preconditoner with layer-wise synchronization strategy
(11) $\tilde{\boldsymbol{\theta}}^{\ast}_{s}:=\operatorname*{arg\,min}_{\tilde{\boldsymbol{\theta}}_{s}\in\mathbb{R}^{n_{s}}}\pazocal{L}(\tilde{\boldsymbol{\theta}}_{s}+(\boldsymbol{\mathrm{I}}-\boldsymbol{R}^{T}_{s}\boldsymbol{R}_{s})\boldsymbol{\theta}),$
@@
(12) $\pazocal{M}_{SL}(\boldsymbol{\theta}):=\boldsymbol{\theta}+\gamma\sum_{s=1}^{N_{sd}}\boldsymbol{R}^{T}_{s}(\tilde{\boldsymbol{\theta}}^{\ast}_{s}-\boldsymbol{R}_{s}\boldsymbol{\theta})=:\boldsymbol{\theta}+\gamma\sum_{s=1}^{N_{sd}}\boldsymbol{d}_{s}=:\boldsymbol{\theta}+\gamma\boldsymbol{d}.$
@@
(13) $\pazocal{L}\left(\boldsymbol{\theta}+\sum_{s=1}^{N_{sd}}\gamma^{*}_{s}\boldsymbol{d}_{s}\right)<\pazocal{L}\left(\boldsymbol{\theta}+\sum_{s=1}^{N_{sd}}\gamma_{s}\boldsymbol{d}_{s}\right),$
for all $\gamma^{*}_{s}\neq\gamma_{s}$.
The set $\{\gamma^{*}_{s}\}_{s=1}^{N_{sd}}$ can be determined by solving the following $N_{sd}$-dimensional minimization problem:
(14) $\min_{\gamma_{1},\ldots,\gamma_{N_{sd}}}\pazocal{L}\left(\boldsymbol{\theta}+\sum_{s=1}^{N_{sd}}\gamma_{s}\boldsymbol{d}_{s}\right).$
@@
for $s=1,\cdots,N_{sd}$. Thus, the values $\{\gamma^{*}_{s}\}_{s=1}^{N_{sd}}$ are determined from the first subdomain to the last one, utilizing a standard, unidimensional, line-search strategy. This guarantees that each search direction $\boldsymbol{d}_{s}$ decreases the value of $\pazocal{L}$ associated with a global network. At the same time, we ensure that the DNN’s bases are constructed by appropriately adjusting the impact of each search direction $\boldsymbol{d}_{s}$ on the global DNN’s approximation. Here, we remark that the search directions $\{\boldsymbol{d}_{s}\}_{s=1}^{N_{sd}}$ are obtained in parallel, while only scalar damping parameters $\{\gamma^{*}_{s}\}_{s=1}^{N_{sd}}$ are obtained sequentially during the subdomain synchronization step.
@@ 3.2 Two-level overlapping additive Schwarz preconditioner
(15) $\boldsymbol{\theta}_{0}=[\boldsymbol{\theta}_{1,1},\ldots,\boldsymbol{\theta}_{s,1},\ldots,\boldsymbol{\theta}_{N_{sd},1}]^{T},$
where $\boldsymbol{\theta}_{s,1}$ is extracted from $\boldsymbol{\theta}_{s}=[\boldsymbol{\theta}_{s,1},\ldots,\boldsymbol{\theta}_{s,\tilde{N}_{s}}]^{T}$, for all $s=1,\ldots,N_{sd}$. An illustration of the coarse-level parameter selection for a six-layer network with $N_{sd}=3$ and $\delta=0$ is depicted in Fig. 3.
The transfer of the data between the coarse-level subnetwork and the global network, is ensured by two transfer operators. In particular, we consider the restriction operator $\boldsymbol{R}_{0}\colon\mathbb{R}^{n}\to\mathbb{R}^{n_{0}}$ and its adjoint, the prolongation operator $\boldsymbol{R}^{T}_{0}\colon\mathbb{R}^{n_{0}}\to\mathbb{R}^{n}$, given such that
$\boldsymbol{\theta}_{0}=\boldsymbol{R}_{0}\boldsymbol{\theta},\qquad\qquad\boldsymbol{\theta}=\boldsymbol{R}_{0}^{T}\boldsymbol{\theta}_{0}.$
Note that the number of trainable coarse-level parameters $n_{0}$ scales proportionally to the number of subdomains $N_{sd}$.
Figure 3: An example of network decomposition associated with coarse level training. The first layer (red color) of each subdomain (red+blue color) is selected to construct a set of trainable coarse level parameters.
@@
(16) $\hat{\boldsymbol{\theta}}=\boldsymbol{\theta}+\boldsymbol{R}^{T}_{0}(\boldsymbol{\theta}^{\ast}_{0}-\boldsymbol{R}_{0}\boldsymbol{\theta}).$
@@
$\pazocal{M}_{TL}(\boldsymbol{\theta}):=\pazocal{M}_{SL}(\underbrace{\boldsymbol{\theta}+\boldsymbol{R}^{T}_{0}(\boldsymbol{\theta}^{\ast}_{0}-\boldsymbol{R}_{0}\boldsymbol{\theta})}_{:=\hat{\boldsymbol{\theta}}}).$
Note that since the coarse-level training indirectly enforces the feed-forward flow of the features through the network, the parameters $\hat{\boldsymbol{\theta}}$ serve as an improved initial guess for the parallel subdomain training phase. Algorithm 1 summarizes the proposed two-level preconditioning process for an user selected optimizer $\pazocal{K}$.
@@
0:  $\pazocal{L}\colon\mathbb{R}^{n}\to\mathbb{R}$, $\boldsymbol{\theta}^{(0)}\in\mathbb{R}^{n}$, $\varepsilon\in\mathbb{R}$, $k=0$, $\delta\in\mathbb{N}$, $\pazocal{K}:\mathbb{R}^{n}\to\mathbb{R}^{n}$
@@
2:     $\boldsymbol{\theta}^{\ast}_{0}=\operatorname*{arg\,min}_{\boldsymbol{\theta}_{0}}\pazocal{L}(\boldsymbol{\theta}_{0}+(\boldsymbol{\mathrm{I}}-\boldsymbol{R}^{T}_{0}\boldsymbol{R}_{0})\boldsymbol{\theta}^{(k)})$ //Coarse level training
@@
4:     for $s=1,\cdots,N_{sd}$ do
5:        $\tilde{\boldsymbol{\theta}}^{\ast}_{s}=\operatorname*{arg\,min}_{\tilde{\boldsymbol{\theta}}_{s}}\pazocal{L}(\tilde{\boldsymbol{\theta}}_{s}+(\boldsymbol{\mathrm{I}}-\boldsymbol{R}^{T}_{s}\boldsymbol{R}_{s})\hat{\boldsymbol{\theta}}^{(k)})$ //Parallel subdomain training
@@
11:     $\boldsymbol{\theta}^{(k+1/2)}=\hat{\boldsymbol{\theta}}^{(k)}+\sum_{s=1}^{N_{sd}}\gamma_{s}\boldsymbol{d}_{s}$ //Subdomain correction recombination step
12:     $\boldsymbol{p}^{(k)}=\pazocal{K}\big{(}\nabla\pazocal{L}(\boldsymbol{\theta}^{(k+1/2)})\big{)}$ //Global optimizer step
13:     $\boldsymbol{\theta}^{(k+1)}=\boldsymbol{\theta}^{(k+1/2)}+\alpha^{(k)}\boldsymbol{p}^{(k)}$ //Global parameter update
@@ 4.1 Benchmark problems for training of PINNs
    $u$ $=-\sin(\pi x),$ $\forall\ (t,x)\in\{0\}\times[-1,1],$
    $u$ $=0,$ $\forall\ (t,x)\in(0,1]\times\{1\},$
    $u$ $=0,$ $\forall\ (t,x)\in(0,1]\times\{-1\},$
@@
    $u$ $=0,$ on $\text{on }\partial\Omega,$
@@
    $u$ $=x^{2}\cos(\pi x),$ $\forall\ (t,x)\in\{0\}\times[-1,1],$
    $u$ $=-1,$ $\forall\ (t,x)\in(0,1]\times\{-1\},$
    $u$ $=-1,$ $\forall\ (t,x)\in(0,1]\times\{1\},$
@@
The amount and location of collocation points significantly influence the discretization error of PINNs. Here, we perform sampling by employing the quasi-Monte Carlo method [57], with Hammersley low-discrepancy sequences [56]. As a consequence, the convergence rate of Quasi-Monte Carlo is close to $O(N^{-1})$, whereas the convergence rate of the standard Monte Carlo method, with pseudorandom sequences, is $O(N^{-1/2})$ [3]. The number of collocation points is $10{,}000$ for all three examples.
Table 1: The summary of network architectures for PINN’s benchmark problems.
@@ 4.2 Benchmark problems for training of DONs
    (17) $-\nabla\cdot(K(x,\boldsymbol{\eta})\ \nabla u(x))$ $=f(x),\quad\quad$ $\forall\ x\in\Omega,$
    $u$ $=0,$ $\forall\ x\in\partial\Omega,$
    where $u$ denotes the solution and $f(x):=1$ is the forcing term. The diffusion coefficient $K(x,\boldsymbol{\eta})$ is considered to be anisotropic tensor of the following form:
@@
    where $\boldsymbol{\eta}:=[\alpha,\beta]$. The parameter ${\beta\in[10^{-6},1]}$ specifies the anisotropic strength, and it is sampled from the following distribution: $\log_{10}(\nicefrac{{1}}{{\beta}})\sim\pazocal{U}[0,6]$. The parameter ${\alpha\in(0,\pi)}$, sampled as $\alpha\sim\pazocal{U}[0,\pi]$, is used to specify the angle of anisotropic direction. The components of the forcing term are selected randomly from the Gaussian distribution $\pazocal{N}(0,1)$. To construct the dataset, for each $\boldsymbol{\alpha}$, we discretize (17) using uniform mesh with $34\times 34$ elements. The number of training samples and test samples are $4{,}250$ and $750$, respectively. We have uploaded this new dataset to the public ZENODO repository [38].
@@
  • Helmholtz equation (Helm): Let $\Omega$ be the computational domain, defined as an annular cylinder depicted on Fig. 4. The parametrized Helmholtz equation is given as
    (18) $-\Delta u(x)-k_{\text{H}}^{2}u(x)$ $=f(x,\boldsymbol{\eta}),\quad\quad\quad$ $\forall\ x\in\Omega,$
    $u$ $=0,$ $\forall\ x\ \text{on}\ \partial\Omega_{T},$
    where $u$ is the solution and $\Omega_{T}$ denotes the top boundary of the cylinder. Moreover, the symbol $k_{\text{H}}$ denotes a constant wave number, which we set to $k_{\text{H}}=3$. We sample the forcing term from Gaussian random fields (GRFs) with mean $\mathbb{E}[f(x,\boldsymbol{\eta})]=0$, and the covariance is given as
@@
    with $\sigma=3.0$ and $\ell=0.1$. To construct the dataset, we discretize (18) using an unstructured mesh with $1,503$ nodes. The sampled right-hand sides are then interpolated to the bounding box $[-1,1]\times[-1,1]\times[0,1]$, consisting of $32\times 32\times 32$ elements, to create a representation that 3D convolutional layers of the branch network can efficiently process. The number of training samples and test samples are $4{,}250$ and $750$, respectively. We have made this new dataset publicly accessible on the ZENODO repository [39].
@@
    $u$ $=u_{0}(x,\boldsymbol{\eta}),$ $\forall\ (x,t)\in\Omega\times\{0\}.$
    The DON is trained to learn a map from the initial condition to the solution of PDE. To this aim, $u_{0}$ is parametrized as $u_{0}(x,\boldsymbol{\eta}):=h_{1}1_{\{c_{1}-\frac{w}{2},c_{1}+\frac{w}{2}\}}\\ +\sqrt{\max(h_{2}^{2}-a^{2}(x-c_{2})^{2},0)}$, where the parameters $\boldsymbol{\eta}:=[a,h_{1},h_{2},c_{1},c_{2},w]$ are sampled randomly from the range $[0.1,0.2]$, $[0.5,2.0]$, $[0.5,2.0]$, $[0.2,0.3]$, $[0.7,0.8]$, for $a$, $h_{1}$, $h_{2}$, $c_{1}$, $c_{2}$ and $w$, respectively. The dataset for training the DON is constructed using a finite difference method with uniform mesh composed of $40\times 40$ nodes in space-time. The number of training and test samples is $1{,}000$. Note that for this benchmark problem, we utilize the POD-DeepONet architecture [53], which replaces the trunk network with the POD bases. This allows us to effectivelly demonstrate the capabilities of the proposed layer-wise decomposition for different types of DON architectures.
Table 2: The summary of network architectures for DON’s benchmark problems. The symbol []delimited-[][$\cdot$][ ⋅ ] indicates the number of channels for CNN and the width for FNN, respectively. The symbols B and T denote the branch network and trunk network, respectively.

### v2

@@ 3.1 Overlapping layer-wise additive Schwarz preconditioner
The computational efficiency associated with solving the preconditioned system of equations (10) is directly linked to the quality and the computational cost of the nonlinear preconditioner $\pazocal{M}$. In this work, we construct the preconditioner $\pazocal{M}$ by utilizing the overlapping layer-wise network decomposition. Thus, we decompose the DNN, consisting of $N$ layers, into $N_{sd}$ subnetworks (subdomains), such that each layer belongs to a unique subnetwork. The parameters of the network are also decomposed into $N_{sd}$ disjoint groups, i.e.,
$\boldsymbol{\theta}=[\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{s},\ldots,\boldsymbol{\theta}_{N_{sd}}]^{T},$
where $\boldsymbol{\theta}_{s}:=[\boldsymbol{\theta}_{s,1},\ldots,\boldsymbol{\theta}_{s,N_{s}}]^{T}\in\mathbb{R}^{n_{s}}$, for $s=1,\ldots,N_{sd}$. Each $s^{th}$ non-overlapping subnetwork consists of $N_{s}$ layers, with parameters of $h^{th}$ layer being denoted as $\boldsymbol{\theta}_{s,h}$. In the case of the overlapping decomposition, the parameters $\boldsymbol{\theta}_{s}$ of the $s$-th subnetwork are concatenated with the parameters of $\delta(\in\mathbb{N})<N_{S}$ neighboring layers from the preceding (s1)(s-1)-th and succeeding $(s+1)$-th subnetworks. For $\delta\geq 1$, this gives rise to overlapping set of subdomain parameters
@@
where $\tilde{N}_{s}$ and $\tilde{n}_{s}$ denotes the number of layers in overlapping subnetwork and the number of associated subnetwork parameters, respectivelly. Note that, for the first ($s=1$) and the last ($s=N_{sd}$-th) subnetworks, we only pick layers from the second and $(N_{sd}-1)$-th subnetworks, respectively. Figure˜2 illustrates the parameter selection for the second subnetwork of the six-layer network decomposed in a layer-wise manner into three subdomains with an overlap of size one, i.e., $N_{sd}=3$ and $\delta=1$.
Remark 1.
The subdomain size can range from a minimum of $1$ (serial) to a maximum of $n$ layers. The overlap size can range from a minimum of 0 (non-overlapping) to the maximum number of layers in a subnetwork ($n_{s}=n/N_{sd}$ if the layers are decomposed uniformly).
@@ 3.1.1 Single-level preconditoner with layer-wise synchronization strategy
$\tilde{\boldsymbol{\theta}}^{\ast}_{s}:=\operatorname*{arg\,min}_{\tilde{\boldsymbol{\theta}}_{s}\in\mathbb{R}^{n_{s}}}\pazocal{L}(\tilde{\boldsymbol{\theta}}_{s}+(\boldsymbol{\mathrm{I}}-\boldsymbol{R}^{T}_{s}\boldsymbol{R}_{s})\boldsymbol{\theta}),$ (11)
@@
$\pazocal{M}_{SL}(\boldsymbol{\theta}):=\boldsymbol{\theta}+\gamma\sum_{s=1}^{N_{sd}}\boldsymbol{R}^{T}_{s}(\tilde{\boldsymbol{\theta}}^{\ast}_{s}-\boldsymbol{R}_{s}\boldsymbol{\theta})=:\boldsymbol{\theta}+\gamma\sum_{s=1}^{N_{sd}}\boldsymbol{d}_{s}=:\boldsymbol{\theta}+\gamma\boldsymbol{d}.$ (12)
@@
$\pazocal{L}\left(\boldsymbol{\theta}+\sum_{s=1}^{N_{sd}}\gamma^{*}_{s}\boldsymbol{d}_{s}\right)<\pazocal{L}\left(\boldsymbol{\theta}+\sum_{s=1}^{N_{sd}}\gamma_{s}\boldsymbol{d}_{s}\right),\quad\forall\gamma^{*}_{s}\neq\gamma_{s},$ (13)
where $\gamma_{s}\in[0,1]$ is an initial damping parameter associated with the local, subdomain-based, search direction $\boldsymbol{d}_{s}$.
The optimal set of parameters $\{\gamma^{*}_{s}\}_{s=1}^{N_{sd}}$ can be determined by solving the following $N_{sd}$-dimensional minimization problem:
$\min_{\gamma_{1},\ldots,\gamma_{N_{sd}}}\pazocal{L}\left(\boldsymbol{\theta}+\sum_{s=1}^{N_{sd}}\gamma_{s}\boldsymbol{d}_{s}\right).$ (14)
@@
for $s=1,\ldots,N_{sd}$. Thus, the values $\{\gamma^{*}_{s}\}_{s=1}^{N_{sd}}$ are determined from the first subdomain to the last one, utilizing a standard, unidimensional, line-search strategy. This guarantees that each search direction $\boldsymbol{d}_{s}$ decreases the value of $\pazocal{L}$ associated with a global network. At the same time, we ensure that the DNN’s bases are constructed by appropriately adjusting the impact of each search direction $\boldsymbol{d}_{s}$ on the global DNN’s approximation. Here, we remark that the search directions $\{\boldsymbol{d}_{s}\}_{s=1}^{N_{sd}}$ are obtained in parallel, while only scalar damping parameters $\{\gamma^{*}_{s}\}_{s=1}^{N_{sd}}$ are obtained sequentially during the subdomain synchronization step.
@@ 3.2 Two-level overlapping additive Schwarz preconditioner
$\boldsymbol{\theta}_{0}=[\boldsymbol{\theta}_{1,1},\ldots,\boldsymbol{\theta}_{s,1},\ldots,\boldsymbol{\theta}_{N_{sd},1}]^{T},$ (15)
where $\boldsymbol{\theta}_{s,1}$ is extracted from $\boldsymbol{\theta}_{s}=[\boldsymbol{\theta}_{s,1},\ldots,\boldsymbol{\theta}_{s,\tilde{N}_{s}}]^{T}$, for all $s=1,\ldots,N_{sd}$. An illustration of the coarse-level parameter selection for a six-layer network with $N_{sd}=3$ and $\delta=0$ is depicted in Figure˜3. The computational cost of the coarse-level problem is proportional to the number of subdomains $N_{sd}$.
The transfer of the data between the coarse-level subnetwork and the global network, is ensured by two transfer operators. In particular, we consider the restriction operator $\boldsymbol{R}_{0}\colon\mathbb{R}^{n}\to\mathbb{R}^{n_{0}}$, given such that
$\boldsymbol{\theta}_{0}=\boldsymbol{R}_{0}\boldsymbol{\theta}.$
The prolongation operator $\boldsymbol{R}^{T}_{0}\colon\mathbb{R}^{n_{0}}\to\mathbb{R}^{n}$ is defined as the adjoint operator of $\boldsymbol{R}_{0}$. Note that the number of trainable coarse-level parameters $n_{0}$ scales proportionally to the number of subdomains $N_{sd}$.
Figure 3: An example of network decomposition associated with coarse level training. The first layer (olive color) of each subdomain (olive+blue color) is selected to construct a set of trainable coarse level parameters.
@@
$\hat{\boldsymbol{\theta}}=\boldsymbol{\theta}+\boldsymbol{R}^{T}_{0}(\boldsymbol{\theta}^{\ast}_{0}-\boldsymbol{R}_{0}\boldsymbol{\theta}).$ (16)
@@
Figure 4: Schematic illustration of the two-level $D$ optimizer for training a DNN decomposed into three subdomains. We first train the coarse-level parameters (shown in olive color), followed by single-level parallel training of the subdomains. The trainable parameters of each subdomain are depicted in different colors, while the non-trainable (fixed) parameters are shown in gray. In the end, the global training step is performed by performing a few iterations of the optimizer $\pazocal{K}$ for all parameters of the DNN.
@@
[…]
Note that since the coarse-level training indirectly enforces the feed-forward flow of the features through the network, the parameters $\hat{\boldsymbol{\theta}}$ serve as an improved initial guess for the parallel subdomain training phase. Figure˜4 provides a graphical illustration of the one-level and two-level Schwarz methods when applied to train the DNNs. In addition, algorithm˜1 summarizes the proposed two-level preconditioning process for a user-selected optimizer $\pazocal{K}$. Our practical implementation of the line-search step (Step 9 in algorithm˜1) employs cubic backtracking with the Strong-Wolfe conditions [59].
@@
[…]
@@
2:  $\boldsymbol{\theta}^{\ast}_{0}=\operatorname*{arg\,min}_{\boldsymbol{\theta}_{0}}\pazocal{L}(\boldsymbol{\theta}_{0}+(\boldsymbol{\mathrm{I}}-\boldsymbol{R}^{T}_{0}\boldsymbol{R}_{0})\boldsymbol{\theta}^{(k)})$ {Coarse level training}
@@
4:  for $s=1,\ldots,N_{sd}$ do
5:   $\tilde{\boldsymbol{\theta}}^{\ast}_{s}=\operatorname*{arg\,min}_{\tilde{\boldsymbol{\theta}}_{s}}\pazocal{L}(\tilde{\boldsymbol{\theta}}_{s}+(\boldsymbol{\mathrm{I}}-\boldsymbol{R}^{T}_{s}\boldsymbol{R}_{s})\hat{\boldsymbol{\theta}}^{(k)})$ {Parallel subdomain training}
@@
11:  $\boldsymbol{\theta}^{(k+1/2)}=\hat{\boldsymbol{\theta}}^{(k)}+\sum_{s=1}^{N_{sd}}\gamma_{s}\boldsymbol{d}_{s}$ {Subdomain correction recombination step}
12:  $\boldsymbol{p}^{(k)}=\pazocal{K}\big(\nabla\pazocal{L}(\boldsymbol{\theta}^{(k+1/2)})\big)$ {Global optimizer step}
13:  $\boldsymbol{\theta}^{(k+1)}=\boldsymbol{\theta}^{(k+1/2)}+\alpha^{(k)}\boldsymbol{p}^{(k)}$ {Global parameter update}
@@ 4.1 Benchmark problems for training of PINNs
    $u$ $=-\sin(\pi x),\$$\forall\ (t,x)\in\{0\}\times[-1,1],$
    $u$ $=0,\$$\forall\ (t,x)\in(0,1]\times\{1\},$
    $u$ $=0,\$$\forall\ (t,x)\in(0,1]\times\{-1\},$
@@
    $u$ $=0,$ $\text{on }\partial\Omega,$
@@
    $u$ $=x^{2}\cos(\pi x),\$$\forall\ (t,x)\in\{0\}\times[-1,1],$
    $u$ $=-1,\$$\forall\ (t,x)\in(0,1]\times\{-1\},$
    $u$ $=-1,\$$\forall\ (t,x)\in(0,1]\times\{1\},$
@@
The amount and location of collocation points significantly influence the discretization error of PINNs. Here, we perform sampling by employing the Quasi-Monte Carlo method [60], with Hammersley low-discrepancy sequences [7]. As a consequence, it is well known that the convergence rate of Quasi-Monte Carlo is close to $O(N^{-1})$, while the convergence rate of the standard Monte Carlo method, with pseudo-random sequences, is $O(N^{-1/2})$ [61]. The number of collocation points is 10,00010{,}000 for all three examples.
Table 1: The summary of network architectures for PINN’s benchmark problems. We use the fully connected neural network (FNN) with hyperbolic tangent (Tanh) activation function for Burg, DA, and AC problems. The symbol $\cdot$] indicates the width for FNN.
@@ 4.2 Benchmark problems for training of DONs
    $-\nabla\cdot(K(x,\boldsymbol{\eta})\ \nabla u(x))$ $=f(x),\quad\quad$ $\forall\ x\in\Omega,$ (17)
    $u$ $=0,\$$\forall\ x\in\partial\Omega,$
    where $u$ denotes the solution and $f(x):=1$ is the forcing term. The diffusion coefficient $K(x,\boldsymbol{\eta})$ is considered to be an anisotropic tensor of the following form:
@@
    where $\boldsymbol{\eta}:=[\alpha,\beta]$. The parameter $\beta\in[10^{-6},1]$} specifies the anisotropic strength, and it is sampled from the following distribution: $\log_{10}(\nicefrac{{1}}{{\beta}})\sim\pazocal{U}[0,6]$}. The parameter $\alpha\in(0,\pi)$}, sampled as $\alpha\sim\pazocal{U}[0,\pi]$, is used to specify the angle of anisotropic direction. To construct the dataset, for each $\boldsymbol{\alpha}$, we discretize (17) using a uniform mesh with $34\times 34$ elements. The number of training samples and test samples are 4,2504{,}250 and $750$, respectively. We have uploaded this new dataset to the public ZENODO repository [63].
@@
  • Helmholtz equation (Helm): Let $\Omega$ be the computational domain, defined as an annular cylinder depicted on Figure˜5. The parametrized Helmholtz equation is given as
    $-\Delta u(x)-k_{\text{H}}^{2}u(x)$ $=f(x,\boldsymbol{\eta}),\quad\quad\quad$ $\forall\ x\in\Omega,$ (18)
    $u$ $=0,\$$\forall\ x\ \text{on}\ \Gamma,$
    $\frac{\partial{u}}{\partial n}$ $=0,\$$\forall\ x\ \text{on}\ \partial\Omega\setminus\Gamma,$
    where $u$ is the solution and $\Gamma$ denotes the top boundary of the cylinder. Moreover, the symbol $k_{\text{H}}$ denotes a constant wave number, which we set to $k_{\text{H}}=3$. We sample the forcing term from Gaussian random fields (GRFs) with mean $\mathbb{E}[f(x,\boldsymbol{\eta})]=0$, and the covariance is given as
@@
    with $\sigma=3.0$ and $\ell=0.1$. To construct the dataset, we discretize (18) using an unstructured mesh with $1,503$ nodes. The sampled right-hand sides are then interpolated onto a $32\times 32\times 32$ grid composed of elements within the bounding box $1,1]\times[-1,1]\times[0,1]$, generating an input that can be processed by the 3D convolutional layers of the branch network. The number of training samples and test samples are 4,2504{,}250 and $750$, respectively. We have made this new dataset publicly accessible on the ZENODO repository [64].
@@
    $u$ $=u_{0}(x,\boldsymbol{\eta}),\$$\forall\ (x,t)\in\Omega\times\{0\}.$
    The DON is trained to learn a map from the initial condition to the solution of the PDE. To this aim, $u_{0}$ is parametrized as
    $u_{0}(x,\boldsymbol{\eta}):=h_{1}1_{\{c_{1}-\frac{w}{2},c_{1}+\frac{w}{2}\}}\\ +\sqrt{\max(h_{2}^{2}-a^{2}(x-c_{2})^{2},0)},$
    where the parameters $\boldsymbol{\eta}:=[a,h_{1},h_{2},c_{1},c_{2},w]$ are randomly sampled from the uniform distributions $a\sim\pazocal{U}(0.1,0.2)$, $h_{1}\sim\pazocal{U}(0.5,2.0)$, $h_{2}\sim\pazocal{U}(0.5,2.0)$, $c_{1}\sim\pazocal{U}(0.2,0.3)$, $c_{2}\sim\pazocal{U}(0.7,0.8)$, and $w\sim\pazocal{U}(0.1,0.2)$, respectively. The dataset for training the DON is constructed using a finite difference method with uniform mesh composed of $40\times 40$ nodes in space-time. The number of training and test samples is 1,0001{,}000. Note that for this benchmark problem, we utilize the POD-DeepONet architecture [48], which replaces the trunk network with the POD bases. This allows us to effectivelly demonstrate the capabilities of the proposed layer-wise decomposition for different types of DON architectures.
Table 2: The summary of network architectures for DON’s benchmark problems (Aniso, Helm and Adv)). The symbol $\cdot$] indicates the number of channels for convolutional neural network (CNN) and the width for fully-connected neural network (FNN), respectively. The symbols B and T denote the branch network and trunk network, respectively. POD-basis is computed by principal component analysis and fixed.